OpenStudy (anonymous):

In parallelogram DEFG, DH = x + 3, HF = 3y, GH = 3x – 3, and HE = 5y + 4. Find the values of x and y.

OpenStudy (anonymous):

Where did the H come from? Missing info?

OpenStudy (anonymous):

its the middle of the parallelogram

OpenStudy (anonymous):

First step is always to draw it out. I'll help you here

OpenStudy (anonymous):



OpenStudy (anonymous):

thats how it look

OpenStudy (anonymous):



OpenStudy (anonymous):

yea !! yours look better

OpenStudy (anonymous):

From the diagram, you can see DH = HF You can form your first equation here

OpenStudy (anonymous):

After that, GH = HE That would be your second equation

OpenStudy (anonymous):

You have two equations and two unknowns. Solve simultaneuosly

OpenStudy (anonymous):

so x+3=3y

OpenStudy (anonymous):

yes. that is your first equation

OpenStudy (anonymous):

whats next

OpenStudy (anonymous):

Find your second equation. I told you how already

OpenStudy (anonymous):

The main thing here is that H is the midpoint of the diagonals. So, half the diagonals would be equal to each other

OpenStudy (anonymous):

H is midpoint of GE. SO GH = HE

OpenStudy (anonymous):

i did that already

OpenStudy (anonymous):

so, what are your two equations?

OpenStudy (anonymous):

x+3=3y and 3x-3=5y+4

OpenStudy (anonymous):

yup. Do you know how to solve simultaneous equations? Using substitution or elimination?

OpenStudy (anonymous):

alittle

OpenStudy (anonymous):

but the y and x together throws me off

OpenStudy (anonymous):

K I'll show you step - by -step

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

x + 3 = 3y --> eqn (1) 3x- 3 = 5y + 4 --> eqn (2)

OpenStudy (anonymous):

From (1), x = 3y - 3 --> (3)

OpenStudy (anonymous):

Substitute x = 3y - 3 into equation (2) 3(3y-3) - 3 = 5y + 4

OpenStudy (anonymous):

Can you solve for y now? Then substitute the value of y into equation 1, 2 or 3 and solve for x

OpenStudy (anonymous):

would i combine like terms?

OpenStudy (anonymous):

3(3y - 3) -3 = 5y + 4 9y - 9 - 3 = 5y + 4 bring y over to one side and the numbers on the other side

OpenStudy (anonymous):

i think i messed up